# Automatically Interpreting Millions of Features in Large Language Models

### v1

@@ 8 Acknowledgements
We would like to acknowledge Lucia Quirke and David Johnston for the work reviewing and discussing the manuscript. We would like to acknowledge Jacob Drori’s role in the initial development of new scoring techniques. We would also like to thank Joseph Bloom, Sam Marks, Can Rager, Jannik Brinkmann and Sam Marks, Neel Nanda, Sarah Schwettmann and Jacob Steinhardt for their discussions and suggestions in an earlier version of the work. We are thankful to Open Philanthropy for funding this work. We are grateful to CoreWeave for providing the compute resources.
@@ Appendix A Appendix
A.1 Explainer Prompt
@@
analyze text and provide an explanation that thoroughly
@@
the text features that are common in the examples, and what
@@
in your explanation.
- Do not make lists of possible explanations.
Keep your explanations short and concise.
@@
explanation, using [EXPLANATION]:
@@
To better find the explanation for the language patterns,
@@
2. Write down general shared features of the text examples.
@@
3. Formulate a hypothesis and write down the final explanation
using [EXPLANATION]:.
@@ 3. Formulate a hypothesis and write down the final explanation
Afterwards an explanation is added to the example {mdframed}
[EXPLANATION]: Common idioms in text conveying positive
@@
A.2 Examples of activating contexts and explanations
As discussed in the main text, the explanations found for a given latent can be very different depending on the way to sample the activating contexts shown to the explainer model, see fig. A1. This has its advantages and disadvantages.
When sampling from only the top activations, it is possible that the explainer model gives an explanation that is more narrow - “The concept of a buffer, referring to something that separates, shields, or protects one thing from another, often used in various contexts such as physical barriers, chemical reactions, or digital data processing” - instead of one that captures the full distribution - “Words or phrases associated with concepts of spatial or temporal separation (buffers, zones, or cushions) or colloquialisms (buff, buffs, or Buffy, referring to a popular TV show or enthusiast)”. Here the narrower explanations resulted in lower scores: the explanation generated from randomly sampled examples scored 0.95 accuracy in fuzzing and 0.93 accuracy in detection, while the explanation generated from top examples only achieves 0.8 accuracy in fuzzing and 0.74 accuracy in detection.
On the other hand, if we look at the second example, sampling from random examples may sometimes confuse the explainer model - “Tokens often precede or succeed prepositions, articles, and words that signal possession or quantity, frequently indicating a relationship between objects or actions. of feature” , which might not see the pattern that is more clear in the top activating examples -“Descriptions of food or events where food is involved, often mentioning leftovers, and sometimes mentioning the act of eating, serving, or storing food, as well as the amount of food, or the pleasure or satisfaction derived from it.”. Here we see very poor scores on the explanation from randomly sampled examples, 0.54 accuracy in detection and 0.57 accuracy in fuzzing, while the explanation generated from the top examples has 0.89 accuracy both in detection and fuzzing (over the full distribution).
Figure A1: Activating contexts of feature 209 (top) and 293 (bottom), from layers 8 and 32, respectively, of the 131k latent SAE trained on the residual stream of Gemma 2 9b. The shown examples are similar to the ones given to the explainer model to come up with explanations.
A.3 Different scoring methods
A.3.1 Detection details
@@
You will be given a certain feature of text, such as
@@
is to determine which examples possess the feature.
@@
Feature explanation: Words related to American football
@@ A.3.2 Fuzzing details
You will be given a certain feature of text, such as ‘‘male
@@
be given a few examples of text that contain this feature.
@@
feature are between tokens << and >>.
@@
Feature explanation: Words related to American football
@@
A.3.3 Surprisal details
In surprisal scoring, the cross-entropy loss of the scorer model is computed over the tokens of an example. For each explanation, this loss is computed with the explanation and with a default explanation - “Various unrelated sentences,” where the examples can either be activating context or non-activating contexts. In our first approach, Llama 3.1 70b base was used. The prompt starts with few shot examples like: {mdframed}
The following is a description of a certain feature of text
and a list of examples that contain the feature.
@@ A.3.3 Surprisal details
 the leading feature which we ought  "
@@
With the pairs of losses with the explanation and with the default explanation, it is possible to compute the decrease in loss caused by having access to the explanation. It is expected that in activating contexts this difference will be greater than in non-activating examples, and the score of the explanation is given by the AUC computed using this loss as a proxy for activating and non-activating labels.
A.3.4 Embedding details
@@
The cosine-similarity between the instruction embedding and the examples is computed and is used as a proxy for activating and non-activating labels when computing the AUC, which is the score of that explanation.
@@ A.3.5 Adversarial examples
Motivated by the phenomenon of feature splitting, we could use “similar” latents to test whether explanations are precise enough to distinguish between similar contexts. A potential approach is using cosine similarity between decoder directions of latents to find counterexamples for an explanation. Some works (Juang et al., 2024; Macgrath, 2024) have shown that a significant fraction of current latent explanations can’t be used to distinguish between features with high similarity.
A.3.6 Correlation between scores
We compute the correlation between these different scores and simulation 600 different latent explanations, of the 131k latent SAE trained on the residual stream of Gemma 2 9b spread across 4 different layers.
Table A1: Pearson correlation computed over 600 different latent scores
@@ A.3.6 Correlation between scores
Fuzzing 1 0.74 0.70 0.42 0.32
Detection 1 0.45 0.70 0.62
Simulation 1 0.34 0.19
@@
A.4 SAE models used
@@
A.5 Factors that influence the explainability
A.5.1 Dependence on dataset
Even though a significant portion of SAE latents being less active when using RPJv2 instead of the Pile, we find that the latents that are active are generally interpretable to the same degree. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.5.1 Dependence on dataset
A.5.2 Dependence on chain of thought and activation information
We find that COT slightly increases the scores of the explanations found, and that it significantly slows down the rate at which one can produce explanations. Giving the explainer model, the activations associated with each token seem to slightly increase the scores of the explanations generated. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.5.2 Dependence on chain of thought and activation information
A.5.3 Dependence on context length
The context length of the shown examples did not significantly change the scores obtained for the explanations. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model were Llama 3.1b 70b instruct, quantized to 4 bits.
@@ A.5.3 Dependence on context length
A.5.4 Dependence on the origin of examples
@@
A.5.5 Dependence on the number of examples
@@
A.5.6 Explainability across layers
@@
Figure A4: Accuracy on fuzzing and detection scoring Dots correspond to the median score over c.a. 300 latent explanations, and colored region denotes the interquartile range.
A.5.7 Dependence on size of explainer and scorer models
Having a larger explainer model leads to better explanations, but our results suggest that it tends to saturate. It is possible that that is due to the fact that the prompt and setup was optimized for Llama 3.1 70b, and that better explanation could be extracted using Claude Sonnet 3.5 if the prompt was better tuned for it. We find that having a smaller model score significantly drops the scores of the explanations—as bad as having the scorer model generate the explanations and the bigger model score them.
Table A7: Comparison of explainer models for Fuzzing and Detection. Numbers shown are the median score and the interquartile (25%-75%) range.
Experiment Fuzzing Detection Embedding
Random explanation 0.51 (0.45–0.57) 0.51 (0.45–0.58) 0.51 (0.44–0.57)
Randomly initialized Topk SAE 0.55 (0.50–0.60) 0.54 (0.50–0.59)
Claude explaining 0.75 (0.68–0.84) 0.75 (0.65–0.85) 0.70 (0.58–0.81)
Llama 70b explaining (4 bit) 0.76 (0.67–0.86) 0.74 (0.63–0.85) 0.67 (0.57–0.80)
Llama 70b explaining (4 bit) 8b scoring 0.69 (0.61–0.77) 0.69 (0.59–0.79)
Llama 8b explaining 0.70 (0.60–0.81) 0.70 (0.59–0.81) 0.64 (0.54–0.75)
Llama 8b explaining, 8b scoring 0.67 (0.59–0.75) 0.67 (0.57–0.77)
Human explaining 0.75 (0.66–0.85) 0.74 (0.64–0.85) 0.71 (0.62–0.81)
A.5.8 Dependence on SAE size and location
@@
A.6 Alignment of latents using decoder weights
See figure A5 for a depiction of the alignment between weights of decoders at various layers. Here, the alignment is measured using the Frobenius inner product between the decoder weight matrices. Due to the permutation symmetry of SAEs layers are not aligned by default, but can’t be aligned using the Hungarian algorithm. We observe that the decoder weights of SAEs trained on the residual stream can be better aligned than those trained on the MLP.
Figure A5: Alignment between decoder directions of the 16k Gemma 2 9b SAEs Because each layer is trained independently, there is no correspondence between latents at different layers. It is possible to align them using the Hungarian algorithm, but even still the MLP SAEs are less aligned than the residual stream ones.
@@ A.7.1 Scoring implementation
For each feature, we sample a pool of 40 length-64 prompts from RPJv2 (Computer, 2023), of which 30 i.i.d. prompts are taken for scoring, while the remaining 10 are used by the explainer. The pool is sampled among nonzero activating contexts, stratified by the quintile of the feature’s max activation in that context (i.e., the nonzero activating documents are sorted by the context’s maximum activation of the current feature, then 8 contexts are sampled from the first quintile, 8 from the second etc.). Each of these prompts is then truncated to include only the first token on which the feature activates, so that all previous activations for the feature are 0. Activations on the first token position are ignored because the SAEs were not trained on that position.
We filter out features that activates on less than 200 of the 10 million RPJ-v2 tokens.
We then sample generations with a maximum of 8 new tokens from the subject model. For generations with the intervention, we perform an additive intervention on the feature at all token positions after and including the final prompt token. We tune the intervention strength of each feature to various KL-divergence values on the scoring set with a binary search. We stop the binary search when the KL divergence is within 10% of the desired value [^4] Sometimes the KL-divergence is not perfectly monotonic in the intervention strength so 10% error is exceeded. We report the average KL divergence that we observe in figure A6. For our zero-ablation experiment, instead of doing an additive intervention we clamp the feature’s activation to 0. Specifically, the hidden states are encoded by the SAE, then the SAE reconstruction error is computed using a clean decoding, then the SAE encoding is clamped and decoded, and the error is added to the clamped decoding.
@@
A.7.2 Generating explanations
We generate explanations using only the intervention’s effect on the subject’s next-token probabilities because this leads to a concise and precise prompt. The explainer, like the scorer, is Llama-3.1 8B.

### v2

@@ 7 Acknowledgements
We would like to acknowledge Lucia Quirke and David Johnston for the work reviewing and discussing the manuscript. We would like to acknowledge Jacob Drori’s role in the initial development of new scoring techniques. We would also like to thank Joseph Bloom, Sam Marks, Can Rager, Jannik Brinkmann and Sam Marks, Neel Nanda, Sarah Schwettmann and Jacob Steinhardt for their suggestions on an earlier version of the work. We are thankful to Open Philanthropy for funding this work. We are grateful to CoreWeave for providing the compute resources.
@@ Appendix A Appendix
A.1 SAE interpretations of a random sentence
In this section, we show some non cherry-picked examples of latent interpretations on a sentence.
Figure A1: SAE latents interpretations for a random sentence. To visualize the latent interpretations produced, we select a sentence taken from the RPJv2 dataset. We selected 4 tokens in different positions in that sentence and filter for latents that are active in different layers. Then we randomly select active latents and their corresponding interpretations to display. We display the detection and fuzzing scores of each interpretation, which indicate how well it explains other examples in the dataset (see Section 3 for details on these scores). The latents selected had high activation, but were not cherry-picked based on interpretations or scores.
A.2 Explainer Prompt
@@
analyze text and provide an interpretation that thoroughly
@@
the text latents that are common in the examples, and what
@@
in your interpretation.
- Do not make lists of possible interpretations.
Keep your interpretations short and concise.
@@
interpretation, using [interpretation]:
@@
To better find the interpretation for the language patterns,
@@
2. Write down general shared latents of the text examples.
@@
3. Formulate a hypothesis and write down the final interpretation
using [interpretation]:.
@@ 3. Formulate a hypothesis and write down the final interpretation
Afterwards an interpretation is added to the example {mdframed}
[interpretation]: Common idioms in text conveying positive
@@
A.3 Examples of activating contexts and interpretations
As discussed in the main text, the interpretations found for a given latent can be very different depending on the way to sample the activating contexts shown to the explainer model, see Figure A2. This has its advantages and disadvantages.
When using the top activating contexts, the explainer model normally gives an interpretation that is more narrow - “The concept of a buffer, referring to something that separates, shields, or protects one thing from another, often used in various contexts such as physical barriers, chemical reactions, or digital data processing” - instead of one that captures the full distribution - “Words or phrases associated with concepts of spatial or temporal separation (buffers, zones, or cushions) or colloquialisms (buff, buffs, or Buffy, referring to a popular TV show or enthusiast)”. Here the narrower interpretations resulted in lower scores: the interpretation generated from examples sampled from the whole distribution scored 0.95 accuracy in fuzzing and 0.93 accuracy in detection, while the interpretation generated from top examples only achieves 0.8 accuracy in fuzzing and 0.74 accuracy in detection.
On the other hand, if we look at the second example, sampling from the whole distribution examples may sometimes confuse the explainer model - “Tokens often precede or succeed prepositions, articles, and words that signal possession or quantity, frequently indicating a relationship between objects or actions. of latent”, which might not see the pattern that is more clear in the top activating examples -“Descriptions of food or events where food is involved, often mentioning leftovers, and sometimes mentioning the act of eating, serving, or storing food, as well as the amount of food, or the pleasure or satisfaction derived from it.” Here we see very poor scores on the interpretation from randomly sampled examples, 0.54 accuracy in detection and 0.57 accuracy in fuzzing, while the interpretation generated from the top examples has 0.89 accuracy both in detection and fuzzing,
Figure A2: Activating contexts of latent 209 (top) and 293 (bottom), from layers 8 and 32, respectively, of the 131k latent SAE trained on the residual stream of Gemma 2 9b. The shown examples are similar to the ones given to the explainer model to come up with interpretations.
A.4 Different scoring methods
A.4.1 Detection details
@@
You will be given a certain latent of text, such as
@@
is to determine which examples possess the latent.
@@
latent interpretation: Words related to American football
@@ A.4.2 Fuzzing details
You will be given a certain latent of text, such as ‘‘male
@@
be given a few examples of text that contain this latent.
@@
latent are between tokens << and >>.
@@
latent interpretation: Words related to American football
@@
A.4.3 Surprisal details
In surprisal scoring, the cross-entropy loss of the scorer model is computed over the tokens of an example. For each interpretation, this loss is computed with the interpretation and with a default interpretation - “Various unrelated sentences,” where the examples can either be activating context or non-activating contexts. In our first approach, Llama 3.1 70b base was used. The prompt starts with few shot examples like: {mdframed}
The following is a description of a certain latent of text
and a list of examples that contain the latent.
@@ A.4.3 Surprisal details
 the leading latent which we ought  "
@@
With the pairs of losses with the interpretation and with the default interpretation, it is possible to compute the decrease in loss caused by having access to the interpretation. It is expected that in activating contexts this difference will be greater than in non-activating examples, and the score of the interpretation is given by the AUC computed using this loss as a proxy for activating and non-activating labels.
A.4.4 Embedding details
@@
The cosine-similarity between the instruction embedding and the examples is computed and is used as a proxy for activating and non-activating labels when computing the AUC, which is the score of that interpretation.
@@ A.4.5 Adversarial examples
Motivated by the phenomenon of latent splitting, we could use “similar” latents to test whether interpretations are precise enough to distinguish between similar contexts. A potential approach is using cosine similarity between decoder directions of latents to find counterexamples for an interpretation. Some works (Juang et al., 2024; Macgrath, 2024) have shown that a significant fraction of current latent interpretations can’t be used to distinguish between latents with high similarity.
A.4.6 Correlation between scores
We compute the correlation between these different scores and simulation 800 different latent interpretations, of the 131k latent SAE trained on the residual stream of Gemma 2 9b spread across 4 different layers. We use 100 activating example, 10 from each of 10 different activating quantiles, and 100 non-activating examples to compute these scores and use Llama 3.1 70b Instruct as the scorer model for simulation, fuzzing and detections, while we use Llama 3.1 70b Base for surprisal and Stella-400M as the embedding model.
Interestingly, we also find that the correlation between simulation scores and other scores also increases when the simulation scorer is Claude Sonnet 3.5. These estimations are on a smaller number of latents(100), due to the high cost of simulation using Claude Sonnet, but on this set, the Spearman correlation between fuzzing and simulation increases from 0.69 to 0.75, between detection and simulation increases from 0.31 to 0.38, between embedding and simulation from 0.14 to 0.21 and from surprisal to simulation from 0.02 to 0.10. The baseline numbers for the correlation of simulation scores are lower on this set, so we expect that had we done Claude scores on a larger set of explanations, the correlations between the scores would be higher.
Table A1: Spearman correlation, computed over scores of 800 different latent interpretations.
@@ A.4.6 Correlation between scores
Fuzzing 1 0.73 0.75 0.41 0.30
Detection 1 0.44 0.71 0.62
Simulation 1 0.28 0.15
Embedding 1 0.79
Surprisal 1
Table A2: Pearson correlation computed over 800 different latent scores
Fuzzing Detection Simulation Embedding Surprisal
Fuzzing 1 0.74 0.74 0.42 0.32
Detection 1 0.46 0.70 0.62
Simulation 1 0.30 0.14
@@
A.4.7 Disagreement between scores
In this section, we will discuss how different types of scores might evaluate explanations differently. In particular, we will focus on an interpretation with fuzzing score but low detection score, high fuzzing score but low simulation score, high simulation score but low embedding score and high embedding score and low simulation score.
For instance, feature 2 of layer 8 of the 131k latent SAE model trained on the residual stream of Gemma 2 9b has a fuzzing score of 0.9 but a detection score of only 0.43. The interpretation given by our pipeline is ”Verbs that link a subject to additional information, often in a formal or descriptive tone.” Most of the activations of this latent are on sentences like ”This $<$is needed$>$ because Magento (…)” or ”For instance, this$<$ is$>$ technically correct syntax”, where <>absent<>< > represent active tokens. In both these examples, the model correctly identifies the highlighted tokens as active during fuzzing, but incorrectly identifies the context as non-active during detection. On the other hand, detection incorrectly identifies sentences like ”pay for things that would prevent larger issues down the road is better in the long run.” or ”Neuroscientist Jack Gallant calls the research a technologic tour de force and says the ultimate decoder would provide vivid” as active. This explanation is too vague on the activating context, leading to a low detection score, but specific enough in the types of tokens that are active, leading to a high fuzz score.
Feature 281 of the same layer has a fuzzing score of 0.97, but a simulation score of only 0.19. Its interpretation is ”URLs or hyperlinks containing query parameters, indicating a request for specific data or actions on a web page.” The simulation score is lower for than the fuzzing score because the model has to decide which parts of the URL the latent is active on, while the fuzzing scorer either is shown links that are highlighted which it will say are active or non-links that are highlighted and are, most likely, not active.
Feature 8 of layer 24 has a simulation score of 0.73, but an embedding score of only 0.43. Its interpretation is ”Prevalence of logical operators and conjunctions in text, including simple addition and conjunction in various contexts, as well as indicators of contrasting or additional information, often used for comparison or to provide supplementary details”. While doing simulation, the model correctly identifies that in the sentence ”the cost of the unit itself, plus installation and construction costs.”, the token ”plus” is active but all the others aren’t. The same is true for sentences like ”Plus my potato bread I made on the ANZAC Day” and ”Cliff Chiang on reintroducing Orion in Wonder Woman! Plus interviews”. On the other hand, these sentences don’t have high embedding similarity with the explanation.
Feature 10 of layer 16 has the opposite situation happening, with an embedding score of 0.74 and a simulation score of only 0.04. Its interpretation is ”Verbs, prepositions, and adverbs that connect clauses or indicate direction, movement, or progression, often in a sequence of actions or events”. In sentences like ”right, three months ¡to close down¿. There was dead silence when the message was read. Everybody waited for Mr. Smith to speak. Mr. Gingham” the simulator model incorrectly identifies ” waited” and the second ” to” as active, and misses the real activation in ”to close down”. The same happens in ”This is ¡going¿ to ¡be¿ one ¡swinging¿ party. Especially if the guests survive Uncle Wilhelm. I love weddings. I love going to them, I love being in”, where the simulator model identified ”love being in” as well as ”love going to” as being active. The embedding model correctly identified that most activating sentences have the mention of motion.
With these examples, we hope to demonstrate that these generated explanations are not perfect, but that we have an easier understanding of their flaws by looking at which cases different scores disagree on.
A.5 SAE models used
@@
A.6 Factors that influence the explainability
A.6.1 Dependence on dataset
Even though a significant portion of SAE latents are less active when using RPJv2 instead of the Pile, we find that the latents that are active are generally interpretable to the same degree. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.6.1 Dependence on dataset
A.6.2 Dependence on chain of thought and activation information
We find that COT slightly increases the scores of the interpretations found, and that it significantly slows down the rate at which one can produce interpretations. Giving the explainer model, the activations associated with each token seem to slightly increase the scores of the interpretations generated. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model where Llama 3.1b 70b instruct, quantized to 4bit.
@@ A.6.2 Dependence on chain of thought and activation information
A.6.3 Dependence on context length
The context length of the shown examples did not significantly change the scores obtained for the interpretations. These evaluations were done using the 131k latent SAE trained on the residual stream of Gemma 2 9b. The scorer and the explainer model were Llama 3.1b 70b instruct, quantized to 4 bits.
@@ A.6.3 Dependence on context length
A.6.4 Dependence on the origin of examples
@@
A.6.5 Dependence on the number of examples
@@
A.6.6 Explainability across layers
@@
Figure A5: Accuracy on fuzzing and detection scoring Dots correspond to the median score over c.a. 300 latent interpretations, and colored region denotes the interquartile range.
A.6.7 Dependence on size of explainer models
Having a larger explainer model leads to better interpretations, but our results suggest that this benefit tends to saturate, as we find that the scores of explanations given by Claude or Llama 70b are very similar. We only did human interpretation of 100 latents. Simulations scores are also only computed over 100 latents.
Table A8: Comparison of scores of interpretations given by different models. Numbers shown are the median score and the interquartile (25%-75%) range.
Experiment Fuzzing Detection Embedding Simulation
Random interpretation 0.51 (0.45–0.57) 0.51 (0.45–0.58) 0.51 (0.44–0.57) -0.02 (-0.02–0.00)
Claude 0.75 (0.68–0.84) 0.75 (0.65–0.85) 0.70 (0.58–0.81) 0.30 (0.28–0.32)
Llama 70b 0.76 (0.67–0.86) 0.74 (0.63–0.85) 0.67 (0.57–0.80) 0.29 (0.28-0.32)
Llama 8b 0.70 (0.60–0.81) 0.70 (0.59–0.81) 0.64 (0.54–0.75) 0.26 (0.23-0.30)
Human 0.75 (0.66–0.85) 0.74 (0.64–0.85) 0.71 (0.62–0.81) 0.36 (0.32–0.39)
A.6.8 Dependence on size of explainer models
We find that both fuzzing, detection and simulation scoring are dependent on the size of the explainer model, that is, given the same explanations, larger models achieve higher scores on average. While Claude and Llama 80b have similar fuzzing and detection scores, simulation, being more complex, is better performed by Claude.
Table A9: Comparison of scorer models for fuzzing, detection and simulation. Numbers shown are the median score and the interquartile (25%-75%) range.
Experiment Fuzzing Detection Simulation
Claude 0.76 (0.66–0.88) 0.72 (0.58–0.87) 0.33 (0.29–0.36)
Llama 70b 0.76 (0.67–0.86) 0.74 (0.63–0.85) 0.29 (0.28–0.32)
Llama 8b 0.70 (0.60–0.81) 0.70 (0.59–0.81) 0.26 (0.25–0.31)
A.6.9 Dependence on SAE size and location
@@ A.7.1 Scoring implementation
For each latent, we sample a pool of 40 length-64 prompts from RPJv2 (Computer, 2023), of which 30 i.i.d. prompts are taken for scoring, while the remaining 10 are used by the explainer. The pool is sampled among nonzero activating contexts, stratified by the quintile of the latent’s max activation in that context (i.e., the nonzero activating documents are sorted by the context’s maximum activation of the current latent, then 8 contexts are sampled from the first quintile, 8 from the second etc.). Each of these prompts is then truncated to include only the first token on which the latent activates, so that all previous activations for the latent are 0. Activations on the first token position are ignored because the SAEs were not trained on that position.
We filter out latents that activates on less than 200 of the 10 million RPJ-v2 tokens.
We then sample generations with a maximum of 8 new tokens from the subject model. For generations with the intervention, we perform an additive intervention on the latent at all token positions after and including the final prompt token. We tune the intervention strength of each latent to various KL-divergence values on the scoring set with a binary search. We stop the binary search when the KL divergence is within 10% of the desired value [^3] Sometimes the KL-divergence is not perfectly monotonic in the intervention strength so 10% error is exceeded. We report the average KL divergence that we observe in Figure A6. For our zero-ablation experiment, instead of doing an additive intervention we clamp the latent’s activation to 0. Specifically, the hidden states are encoded by the SAE, then the SAE reconstruction error is computed using a clean decoding, then the SAE encoding is clamped and decoded, and the error is added to the clamped decoding.
@@
A.7.2 Generating interpretations
We generate interpretations using only the intervention’s effect on the subject’s next-token probabilities because this leads to a concise and precise prompt. The explainer, like the scorer, is Llama-3.1 8B.